OpenStudy (anonymous):

f(x)=x^(3x) Use logarithmic differentiation to determine the derivative. f'(x)= f'(1)=

OpenStudy (agent0smith):

Change the f(x) to y for convenience, take ln of both sides: \[\Large \ln y= \ln x^{3x}\] then bring down the 3x \[\Large \ln y= 3x \ln x\] Now use implicit differentiation - derivative of ln y, but don't forget the derivative of y is dy/dx then use the product rule like last time on the right\\[\Large \left( \frac{ 1 }{ y } \right)\frac{ dy }{ dx } = 3 \ln x + 3x \left( \frac{ 1 }{ x } \right)\]

OpenStudy (anonymous):

so then i solve for dy/dx

OpenStudy (agent0smith):

Yes, and remember that the y in that equation can be replaced by x^(3x)

OpenStudy (anonymous):

alright give me a minute to try it

OpenStudy (anonymous):

just to make sure i can rewite it as this right [3lnx+3x/x]y which it would end up as [3lnx+3x/x](x^3x)

OpenStudy (agent0smith):

Yeah and you can simplify the 3x/x \[\LARGE \frac{ dy }{ dx } = x^{3x}\left( 3 \ln x + 3 \right)\]

OpenStudy (anonymous):

oh ok then from there i just plug in 1 right

OpenStudy (agent0smith):

Yep :)